# Joint Score-Threshold Optimization for Interpretable Risk Assessment Under Partial Supervision

### v1

@@ 2.4 Governance and Interpretability Constraints
Sign constraints encode clinical knowledge about risk factor directionality. For factors known to increase risk (e.g., age $\geq$ 80, fall history), we enforce $\beta_{j}\geq 0$. For protective factors (e.g., independent mobility), we enforce $\beta_{j}\leq 0$. This prevents counterintuitive weights that would undermine clinical trust:
$\beta_{j}\geq 0\quad\forall j\in\mathcal{J}_{\text{risk}},\quad\beta_{j}\leq 0\quad\forall j\in\mathcal{J}_{\text{protective}}.$ (5)
@@
Sparsity constraints limit cognitive load by bounding the number of active features. Clinical tools typically use 5-10 items for practical usability:
@@
where binary variables $u_{j}$ indicate whether feature $j$ is active. The constraint ensures that at most $s_{\max}$ features have non-zero weights.
Minimal modification constraints facilitate adoption by limiting changes from incumbent tools. Clinicians are more likely to accept tools that refine rather than replace existing practice:
$|\beta_{j}-\beta_{j}^{(0)}|\leq\Delta_{j}\quad\text{or penalize}\quad\rho\sum_{j=1}^{p}|\beta_{j}-\beta_{j}^{(0)}|,$ (7)
where $\beta^{(0)}$ represents the incumbent weights. The first form enforces hard limits on individual changes, while the second adds a soft penalty to the objective.
Grouping constraints ensure related features are treated consistently. For example, different mobility assessments might be grouped:
$\beta_{j}=\beta_{j^{\prime}}\quad\forall(j,j^{\prime})\in\mathcal{G},$ (8)
where $\mathcal{G}$ contains pairs of features that should have identical weights.
Performance constraints ensure the optimized model meets minimum clinical safety requirements. Using the assignment variables $z_{ik}$ from the MIP formulation, we can directly bound error rates. For instances with known true labels (singleton feasible sets $\mathcal{S}_{i}=\{k_{i}^{*}\}$), we define:
False positive rate constraint (e.g., limiting low-risk patients classified as high-risk):
$\frac{\sum_{i:k_{i}^{*}=1}z_{iK}}{|\{i:k_{i}^{*}=1\}|}\leq\alpha_{\text{FP}}.$ (9)
False negative rate constraint (e.g., limiting high-risk patients classified as low-risk):
$\frac{\sum_{i:k_{i}^{*}=K}z_{i1}}{|\{i:k_{i}^{*}=K\}|}\leq\alpha_{\text{FN}}.$ (10)
@@
More generally, for any true category $k^{*}$ and predicted category $k$, we can constrain:
$\sum_{i\in\mathcal{I}_{k^{*}}}z_{ik}\leq\alpha_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|\quad\text{or}\quad\geq\gamma_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|,$ (11)
@@
where $\mathcal{I}_{k^{*}}=\{i:\mathcal{S}_{i}=\{k^{*}\}\}$ is the set of instances with true label $k^{*}$, and $\alpha_{k^{*},k}$ (or $\gamma_{k^{*},k}$) specifies the maximum (or minimum) acceptable rate.
@@
These constraints are linear in the assignment variables $z_{ik}$ and integrate directly into the MIP formulation (19). They are particularly valuable for ensuring that optimization does not sacrifice critical safety metrics (e.g., sensitivity for high-risk patients) in favor of overall accuracy. Note that these constraints only apply to instances with known ground truth (singleton feasible sets), which aligns with the partial supervision framework.
@@ 2.5 Constrained Score Optimization (CSO) Relaxation
The MIP formulation provides exact solutions but faces computational challenges for large datasets. We derive a convex relaxation that preserves the essential structure—ordinal relationships, partial supervision, and asymmetric costs—while enabling efficient optimization via gradient-based methods.
The MIP model uses binary variables to encode category assignments and threshold crossings. The CSO relaxation replaces these discrete decisions with continuous margin-based losses. An important insight is that the ordinal loss decomposes into a sum of boundary crossing penalties, which we can approximate smoothly. Consider predicting category $k$ for an instance with true category $k^{*}$. The ordinal loss $|k-k^{*}|$ equals the number of thresholds between $k$ and $k^{*}$. We can rewrite this as:
@@
$\phi_{i,k}^{-}$ $=\frac{1}{\alpha}\log\left(1+\exp(\alpha(\tau_{k}-s_{i}))\right)$ (12)
[…]
Similarly, $\phi_{i,k}^{+}=\frac{1}{\alpha}\log(1+\exp(\alpha(s_{i}-\tau_{k})))$ penalizes exceeding threshold $k$ when we shouldn’t. The temperature parameter $\alpha$ controls the approximation quality: as $\alpha\to\infty$, softplus approaches the hinge loss; as $\alpha\to 0$, it becomes linear.
To derive the CSO objective for instance $i$ with feasible set $\mathcal{S}_{i}$, we define the CSO loss function as:
[…]
The weights $\lambda_{j}^{+},\lambda_{j}^{-}>0$ encode the asymmetric importance of each boundary. To reproduce the asymmetric ordinal loss (3), we set:
$\lambda_{j}^{-}$ $=\alpha_{\text{under}}\cdot w_{j}\quad\text{(weight for failing to exceed threshold }j\text{)},$ (14)
$\lambda_{j}^{+}$ $=\alpha_{\text{over}}\cdot w_{j}\quad\text{(weight for incorrectly exceeding threshold }j\text{)}.$ (15)
where $w_{j}$ encodes the positional importance (e.g., $w_{j}=1$ for linear growth, $w_{j}=j$ for quadratic).
The minimization over $k\in\mathcal{S}_{i}$ implements partial supervision: the loss encourages the instance toward the most compatible category within its feasible set. This is convex since it’s the minimum of convex functions over a finite set.
The full CSO formulation can be written as follows.
@@
$\tau_{\text{High}}-\tau_{\text{Low}}\geq\Delta_{\min}\quad\text{(for $K=3$)}$ (16d)
$\beta\in\Omega$ (16e)
@@
The CSO objective is differentiable with respect to both $\beta$ and $\tau$, with gradients
$\frac{\partial\mathcal{L}_{i}}{\partial\beta}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\left[\sum_{j=1}^{k-1}\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))-\sum_{j=k}^{K-1}\lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))\right]x_{i},$ (17)
$\frac{\partial\mathcal{L}_{i}}{\partial\tau_{j}}$ $=\sum_{k\in\mathcal{S}_{i}^{*}}p_{ik}\begin{cases}-\lambda_{j}^{+}\sigma(\alpha(s_{i}-\tau_{j}))&\text{if }j<k\\ \lambda_{j}^{-}\sigma(\alpha(\tau_{j}-s_{i}))&\text{if }j\geq k\end{cases},$ (18)
where $\sigma(t)=1/(1+\exp(-t))$ is the sigmoid function, and $\mathcal{S}_{i}^{*}\subseteq\mathcal{S}_{i}$ contains the categories achieving the minimum in (13). When $|\mathcal{S}_{i}^{*}|>1$, we use subgradients or smooth approximations. The projection operators maintain feasibility:
  • $\Pi_{\Omega}(\beta)$: projects onto governance constraints (e.g., sign, sparsity)
  • $\Pi_{\mathcal{T}}(\tau)$: projects onto ordered thresholds with minimum gaps via isotonic regression with spacing constraints
Proposition 1 (Convexity).
The CSO objective (16) is convex in $(\beta,\tau)$ jointly.
Proof sketch.
The softplus functions $\phi_{i,k}^{\pm}$ are convex. The weighted sum in (13) preserves convexity. The minimum over $k\in\mathcal{S}_{i}$ is convex as the pointwise minimum of convex functions. The constraints form a convex set. ∎
Proposition 2 (Approximation Quality).
As $\alpha\to\infty$, the CSO solution approaches the MIP solution on separable data.
Proposition 2 follows from the softplus converging to the hinge loss, which exactly captures threshold crossing violations.
Algorithm 1 CSO-Guided Mixed-Integer Optimization for Risk Scoring
1:Training data $\{(x_{i},\mathcal{S}_{i},w_{i})\}_{i=1}^{n}$, parameters $\delta,\Delta_{\min},\alpha$, convergence tolerance $\epsilon$
2:Scoring weights $\beta^{*}$, thresholds $\tau^{*}$
3:Phase 1: CSO Warm-Start
4:Initialize $\beta^{(0)},\tau^{(0)}$ from incumbent tool or random
5:for $t=1,2,\ldots$ until convergence do
6:  Compute CSO loss gradients via (13)
7:  $\beta^{(t+1)}\leftarrow\Pi_{\Omega}(\beta^{(t)}-\eta_{t}\nabla_{\beta}\mathcal{L})$
8:  $\tau^{(t+1)}\leftarrow\Pi_{\mathcal{T}}(\tau^{(t)}-\eta_{t}\nabla_{\tau}\mathcal{L})$ $\triangleright$ Maintain ordering and gaps
9:  if $\|\beta^{(t+1)}-\beta^{(t)}\|+\|\tau^{(t+1)}-\tau^{(t)}\|<\epsilon$ then
10:   break
11:  end if
12:end for
[…]
14:Phase 2: MIP Refinement
15:Initialize MIP solver with warm-start solution $(\beta_{\text{warm}},\tau_{\text{warm}})$
16:Set initial bounds: $\beta_{j}\in[\beta_{\text{warm},j}-\rho,\beta_{\text{warm},j}+\rho]$
17:Solve MIP (19) with warm-start and bounds
18:Optional: If integer scores required, add $\beta_{j}\in\mathbb{Z}$ constraints
19:return optimal solution $(\beta^{*},\tau^{*})$
@@ 2.6 Two-Phase Optimization: CSO Warm-Start with MIP Refinement
Algorithm 1 presents our two-phase optimization approach. The CSO relaxation in Phase 1 provides a high-quality warm-start solution that significantly accelerates the subsequent MIP solve in Phase 2. This hybrid approach combines the computational efficiency of convex optimization with the exactness of mixed-integer programming. The warm-start not only reduces solve time but also helps avoid poor local optima by providing the MIP solver with a good initial feasible region.
While our formulation allows continuous-valued scoring weights $\beta\in\mathbb{R}^{p}$, practical deployment often benefits from integer-valued scores that are easier for clinicians to compute mentally. Following the approach in [25], we can enforce integrality by adding constraints $\beta_{j}\in\mathbb{Z}$ to the MIP formulation. When combined with appropriate scaling of features, this yields point values similar to existing tools (e.g., 0-5 points per item). The trade-off between interpretability and predictive performance can be explicitly controlled through the granularity of allowed integer values. Our CSO warm-start remains valuable even with integer constraints, as rounding the continuous solution provides an excellent initial integer feasible solution.
@@ 3 Experiments
The Johns Hopkins Fall Risk Assessment Tool provides an ideal test case for our framework as it exemplifies all the key challenges we address: partial supervision due to intervention-censored outcomes, asymmetric costs where missing a fall is more harmful than over-treatment, and the need for interpretable integer-valued scores that clinicians can compute manually. In mapping JHFRAT to our general formulation, the feature vector $x_{i}\in\mathbb{R}^{17}$ represents the 17 binary risk factors across seven clinical categories (age, fall history, mobility, etc.), where $x_{ij}=1$ indicates the presence of risk factor $j$ for patient $i$. The scoring function $s(x_{i})=\beta^{\top}x_{i}$ computes the total JHFRAT score, where $\beta\in\mathbb{Z}^{17}_{+}$ contains the integer point values for each risk factor.
The partial supervision challenge manifests clearly in this setting: among 54,209 hospital encounters, the vast majority (80.7%) received targeted fall prevention interventions that obscure their true risk. Following our framework, we construct feasible label sets $\mathcal{S}_{i}$ based on observed outcomes and intervention patterns. Patients who fell despite interventions are confidently labeled as high-risk ($\mathcal{S}_{i}=\{3\}$), while those who remained fall-free without targeted interventions are labeled as low-risk ($\mathcal{S}_{i}=\{1\}$). The remaining patients who received interventions but did not fall have uncertain risk—they might be truly high-risk patients successfully protected by interventions, or lower-risk patients who received unnecessary precautions. Rather than forcing labels on these ambiguous cases, our framework excludes them from the primary optimization while using them for post-hoc validation.
We evaluate our framework on optimizing the Johns Hopkins Fall Risk Assessment Tool using electronic health record data from three Johns Hopkins Health System hospitals: Johns Hopkins Hospital, Johns Hopkins Bayview Medical Center, and Howard County Medical Center. Our dataset comprises 54,209 hospital admissions between March 28, 2022 and October 27, 2023, with complete JHFRAT assessments, intervention records, and fall events. The JHFRAT assessment consists of 17 non-zero coefficient risk factors across seven categories: age, bladder/bowel elimination, cognition, fall history, patient care equipment, medications, and mobility. On average, patients are assessed twice daily with JHFRAT, and our dataset has an average of 1.87 JHFRAT records per patient day.
@@ 3.1 Dataset Construction and Partial Labels
A total of 498 hospital encounters in the dataset include at least one fall event, constituting 0.92% of encounters and equating to 1.07 falls per 1,000 patient-days. The rarity of fall events, combined with widespread use of preventive interventions, creates the selective labeling challenge central to our approach. The majority of patients in the data receive one or more fall-prevention interventions during hospitalization. Of the 31 possible fall-prevention interventions, 13 were identified by a team of clinicians as “targeted” for being resource-intensive and capable of meaningfully altering fall risk. Such targeted interventions include the use of bed exit alarms, increased patient rounding, and constant monitoring. We assume that these targeted interventions, due to their intensive nature and established efficacy, can meaningfully obscure underlying fall risk. In contrast, we do not consider general interventions, like education to patients and family about fall risks, to be risk-obscuring as they are universally applied and have minimal direct impact on fall occurrence.
Figure 1 illustrates how we partition the 54,209 patient encounters into three cohorts based on fall events and targeted intervention receipt. The key insight is that intervention receipt creates asymmetric information about true risk: while falls indicate high risk regardless of interventions (since interventions failed to prevent them), the absence of falls has different interpretations depending on intervention status. We partition patient encounters into three cohorts with distinct labeling strategies:
  • Fall cohort: All encounters with at least one fall event, regardless of targeted interventions applied. These encounters are labeled Safe High-Risk.
  • Reference cohort: Encounters with no targeted interventions throughout the encounter, and no fall events. These encounters are labeled Safe Low-Risk.
@@
The extreme imbalance in cohort sizes–with 80.7% of patients in the ambiguous intervention cohort–underscores why standard supervised learning approaches fail in this setting. Simply assigning these 43,799 patients to a risk category based on their JHFRAT scores or fall outcomes would introduce massive label noise, as we would be guessing the counterfactual (what would have happened without interventions) for the majority of our training data. Our partial supervision approach instead acknowledges this uncertainty explicitly, training only on the 10,410 patients (19.3%) for whom we have confident labels while reserving the intervention cohort for post-optimization validation to assess how the optimized models generalize to the ambiguous majority.

### v2

@@ 3.3 Governance Constraints
The feasible set $\Omega$ in (1g) encodes interpretability and deployment requirements essential for clinical adoption. These constraints ensure that optimized models remain transparent, clinically valid, and aligned with existing practice. Clinical knowledge about risk factor directionality can be encoded through sign restrictions on the weights. For factors known to increase risk, we enforce $\beta_{j}\geq 0$, while we enforce $\beta_{j}\leq 0$ for protective factors. This prevents counterintuitive weights that would undermine clinical trust, $\beta_{j}\geq 0\quad\forall j\in\mathcal{J}_{\text{risk}},$ and $\beta_{j}\leq 0\quad\forall j\in\mathcal{J}_{\text{protective}}.$
@@
To limit cognitive load during manual assessment, we can bound the number of active features through sparsity constraints. Clinical tools typically use 5–10 items for practical usability. Introducing binary variables $u_{j}$ to indicate whether feature $j$ is active, we enforce:
@@
where $s_{\max}$ bounds the total number of features with non-zero weights.
Clinical adoption is often facilitated by limiting changes from incumbent tools, as clinicians are more likely to accept refinements rather than wholesale replacements of existing practice. Let $\beta^{(0)}$ represent the incumbent weights. We can either enforce hard limits on individual changes or add a soft penalty to the objective:
$|\beta_{j}-\beta_{j}^{(0)}|\leq\Delta_{j}\quad\text{or penalize}\quad\rho\sum_{j=1}^{p}|\beta_{j}-\beta_{j}^{(0)}|.$ (4)
The first form constrains each weight individually, while the second adds an $L_{1}$ penalty that encourages overall similarity to the incumbent model.
When features are inherently ordinal, monotonicity constraints can enforce the proper ordinal relationship. For example, the risk coefficient for taking two high fall risk medications should be at least as large as the coefficient for taking one high fall risk medication. Let $\mathcal{G}$ denote the set of feature pairs that should be ordered: $\beta_{j}\leq\beta_{j^{\prime}}\quad\forall(j,j^{\prime})\in\mathcal{G}.$
All of these interpretability constraints can be made to be linear and thus integrated directly into the MIP formulation, allowing optimization solvers to efficiently navigate the constrained solution space while maintaining clinical validity. The specific constraints deployed depend on the clinical context. Sign restrictions are nearly universal, sparsity is important when manual computation is required, minimal modification eases adoption of updated tools, and ordering ensures consistency and robustness against training data noise when features measure related constructs.
3.4 Performance Constraints
Performance constraints, encoded via the domain constraints in (1h), ensure the optimized model meets minimum safety and efficacy requirements. Using the assignment variables $z_{ik}$ from the MIP formulation, we can directly constrain classification error rates and bound operational metrics like sensitivity and specificity. For false positive rate control, we can limit the fraction of low-risk patients misclassified into higher risk categories. For example, to ensure that no more than $\xi_{\text{FPR}}$ fraction of truly low-risk patients are classified as high-risk:
$\frac{\sum_{i:k_{i}^{*}=1}z_{iK}}{|\{i:k_{i}^{*}=1\}|}\leq\xi_{\text{FPR}}.$ (5)
Similarly, for false negative rate control, we can limit the fraction of high-risk patients misclassified as low-risk. To ensure sensitivity of at least $1-\xi_{\text{FNR}}$ for high-risk detection:
$\frac{\sum_{i:k_{i}^{*}=K}z_{i1}}{|\{i:k_{i}^{*}=K\}|}\leq\xi_{\text{FNR}}.$ (6)
@@ 3.4 Performance Constraints
More generally, for any true category $k^{*}$ and predicted category $k$, we can impose upper or lower bounds on misclassification rates:
$\sum_{i\in\mathcal{I}_{k^{*}}}z_{ik}\leq u_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|\quad\text{or}\quad\geq l_{k^{*},k}\cdot|\mathcal{I}_{k^{*}}|,$ (7)
@@
where $\mathcal{I}_{k^{*}}=\{i:\mathcal{S}_{i}=\{k^{*}\}\}$ denotes instances with known true label $k^{*}$, and $u_{k^{*},k}$ (or $l_{k^{*},k}$) specifies the maximum (or minimum) acceptable rate.
@@
These constraints are linear in the assignment variables and integrate directly into the MIP formulation (1). They are particularly valuable for ensuring that optimization does not sacrifice critical safety metrics (e.g., sensitivity for high-risk patients) in favor of overall accuracy. In practice, performance constraints can be derived from incumbent tool performance (ensuring the new model does not regress on key metrics) or from external regulatory or quality requirements.
@@ 3.5 Constrained Score Optimization (CSO) Relaxation
The MIP formulation provides exact solutions but faces computational challenges for large datasets. We derive a convex relaxation that preserves the essential structures of ordinal relationships, partial supervision, and asymmetric costs while enabling efficient optimization via gradient-based methods. The MIP model uses binary variables to encode category assignments and threshold crossings. The CSO relaxation replaces these discrete decisions with continuous margin-based losses. An important insight is that the ordinal loss decomposes into a sum of boundary crossing penalties, which we can approximate smoothly.
Consider predicting category $k$ for an instance with true category $k^{*}$. The ordinal loss $|k-k^{*}|$ equals the number of thresholds between $k$ and $k^{*}$. We can rewrite this as:
@@
$\phi_{i,k}^{-}$ $=\log\left(1+\exp(\alpha^{-}(\tau_{k}-s_{i}))\right)$ (8)
$\approx\begin{cases}0&\text{if }s_{i}\gg\tau_{k}\\ \alpha^{-}(\tau_{k}-s_{i})&\text{if }s_{i}\ll\tau_{k}\\ \log(1+\exp(\alpha^{-}))&\text{if }s_{i}=\tau_{k}\end{cases}.$
Similarly, $\phi_{i,k}^{+}=\log(1+\exp(\alpha^{+}(s_{i}-\tau_{k})))$ penalizes exceeding threshold $k$ when we should not. The temperature parameters $\alpha^{-}$ and $\alpha^{+}$ control the steepness of the linear penalty for the score falling on the wrong side of the threshold.
To derive the CSO objective for instance $i$ with category label $k_{i}^{*}$, we define the CSO loss function as:
$\mathcal{L}_{i}(\beta,\tau)=\min_{k\in\mathcal{S}_{i}}\sum_{j=1}^{k_{i}^{*}-1}\lambda_{j}^{+}\phi_{i,j}^{+}+\sum_{j=k_{i}^{*}}^{K-1}\lambda_{j}^{-}\phi_{i,j}^{-},$ (9)
where the weights $\lambda_{j}^{+},\lambda_{j}^{-}>0$ encode the asymmetric importance of each boundary. The full CSO formulation can be written as follows:
@@
$\beta\in\Omega$ (10d)
@@
The CSO objective is smooth and thus differentiable everywhere with respect to both $\beta$ and $\tau$. Furthermore, the CSO objective (10) is convex in $(\beta,\tau)$ jointly because the softplus functions $\phi_{i,k}^{\pm}$ are convex. The weighted sum in (9) preserves convexity. The differentiability and convexity of the objective, along with the convexity of the feasible set formed by the constraints allows for efficient solution of CSO using commercial nonlinear solvers.
@@ 3.6 Two-Phase Optimization: CSO Warm-Start with MIP Refinement
The CSO relaxation can provide a high-quality warm-start solution that significantly accelerates the subsequent MIP solve. The major challenge with implementing this warm-start approach is ensuring MIP feasibility of the warm-start solution. Simple rounding can be applied to adhere to integrality constraints for the risk model coefficients. Any continuous linear governance constraints from the MIP can be added to the CSO formulation without compromising differentiability and convexity. Performance constraints, on the other hand, cannot be directly added to CSO due to their reliance on the MIP binary variables. However, hyperparameter tuning of the CSO $\alpha$ and $\lambda$ values can be used to find solutions which satisfy these constraints.
This hybrid approach of using CSO and MIP combines the computational efficiency of convex optimization with the exactness of mixed-integer programming. The warm-start reduces solve time and helps avoid poor local optima by providing the MIP solver with a good initial feasible region.
@@ 4 Experiments
The Johns Hopkins Fall Risk Assessment Tool provides an ideal test case for our framework as it exemplifies all the key challenges we address: labels available only for extreme categories due to intervention-censored outcomes, asymmetric costs where missing a fall is more harmful than over-treatment, and the need for interpretable integer-valued scores that clinicians can compute manually. In mapping JHFRAT to our general formulation, the feature vector $x_{i}\in\{0,1\}^{17}$ represents the 17 binary risk factors across seven clinical categories (age, fall history, mobility, etc.), where $x_{ij}=1$ indicates the presence of risk factor $j$ for patient $i$. The scoring function $s(x_{i})=\beta^{\top}x_{i}$ computes the total JHFRAT score, where $\beta\in\mathbb{Z}^{17}_{+}$ contains the integer point values for each risk factor.
We evaluate our framework on optimizing the Johns Hopkins Fall Risk Assessment Tool using electronic health record data from three Johns Hopkins Health System hospitals: Johns Hopkins Hospital, Johns Hopkins Bayview Medical Center, and Howard County Medical Center. Our dataset comprises 54,209 hospital admissions between March 28, 2022 and October 27, 2023, with complete JHFRAT assessments, intervention records, and fall events. The JHFRAT assessment consists of 17 non-zero coefficient risk factors across seven categories: age, bladder/bowel elimination, cognition, fall history, patient care equipment, medications, and mobility. On average, patients are assessed twice daily with JHFRAT, and our dataset has an average of 1.87 JHFRAT records per patient day. This study was approved by the Johns Hopkins University Institutional Review Board: IRB00395329, Functional measures in the hospital and patient outcomes.
@@ 4.1 Dataset Construction and Partial Labels
A total of 498 hospital encounters in the dataset include at least one fall event, constituting 0.92% of encounters and equating to 1.07 falls per 1,000 patient-days. The majority of patients in the data receive one or more fall-prevention interventions during hospitalization. Of the 31 possible fall-prevention interventions, 13 were identified by a team of clinicians as “targeted” for being resource-intensive and capable of meaningfully altering fall risk. Such targeted interventions include the use of bed exit alarms, increased patient rounding, and constant monitoring. We assume that these targeted interventions, due to their intensive nature and established efficacy, can meaningfully obscure underlying fall risk. In contrast, we do not consider general interventions, like educating patients to wait/call for assistance prior to mobilizing, to be risk-obscuring as they are universally applied and have minimal direct impact on fall occurrence. The partial supervision challenge manifests clearly in this setting: among 54,209 hospital encounters, the vast majority (80.7%) received targeted fall prevention interventions that obscure their true risk. The rarity of fall events, combined with widespread use of preventive interventions, creates the selective labeling challenge central to our approach.
For patients who experienced any falls, we consider only the JHFRAT assessments prior to the first fall event. To obtain a medically representative sample for the study cohort, we include patients who fell between the 2nd and 21st days of the encounter, or whose overall length of stay was between 2 and 21 full days. We exclude all encounters with fewer than two JHFRAT assessments during the included encounter time. After applying this filtering, the study cohort includes 46,991 patient encounters.
Figure 1 illustrates how we partition the 46,991 filtered patient encounters into three cohorts based on fall events and targeted intervention receipt. The key insight is that intervention receipt creates asymmetric information about true risk: while falls indicate high risk regardless of interventions (since interventions failed to prevent them), the absence of falls can only be construed as an indicator of true low risk if the patient received no targeted interventions.
We partition patient encounters into three cohorts with distinct labeling strategies:
  • Fall cohort: All encounters with at least one fall event, regardless of targeted interventions applied. These encounters are labeled high-risk.
  • Reference cohort: Encounters with no targeted interventions throughout the encounter, and no fall events. These encounters are labeled low-risk.
@@
The extreme imbalance in cohort sizes underscores why standard supervised learning approaches fail in this setting. Simply assigning the 38,643 intervention cohort patients to a risk category based on their JHFRAT scores or fall outcomes would introduce massive label noise, as we would be guessing the counterfactual (what would have happened without interventions) for the majority of our training data. Our partial supervision approach instead acknowledges this uncertainty explicitly, training only on the 8,348 patients (17.8% of included encounters) for whom we have confident labels while reserving the intervention cohort for post-optimization validation to assess how the optimized models generalize to the ambiguous majority.